OpenStudy (anonymous):

Chris, who is 6 feet tall, is walking away from a street light pole 30 feet high at a rate of 2 feet per second. How fast is the tip of his shadow moving when he is 24 feet from the pole? @Physics

OpenStudy (tiff):

Similar to this problem. x3.9 # 13. A street light is mounted at the top of a 15-ft-tall pole. A man 6 ft tall walks away from the pole with a speed of 5 ft/s along a straight path. How fast is the tip of his shadow moving when he is 40 ft from the pole? Solution. Use similar triangles. Let d be the distance between the man and the pole and let x be the length of the man's shadow. Then x=(d + x) = 6=15, so 15x = 6x + 6d and 9x = 6d. Then x = (2=3)d. The problem wants us to nd d(d + x)=dt = d(d)=dx + dx=dt, since this is the speed of the tip of his shadow. We know dx=dt = (2=3)d(d)=dt and we know d(d)=dt = 5 ft/s, from the problem. Then d(d + x)=dt = (5=3)d(d)=dt = 25=3 ft/s.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you stink